x + y = 4k
2x = k
where k is a constant Real Number, solve the system for (x; y) in terms of k. Give your
solution as an ordered pair.

Answer by ewatrrr(24785)

   
Hi,
x + y = 4k
2x = k or x = k/2
k/2 + y = 4k = (8/2)k
y = (7/2)k
Ordered Pair (k/2, (7/2)k)